# External charged debris in a flowing plasma : charge fluctuation induced complexity

### v2

@@ Abstract
[…]
organization=Physics Department, Gauhati University,city=Guwahati, postcode=781014, state=Assam, country=India
@@
Model: Investigates the effect of periodic charge fluctuation of external debris, induced by ion-acoustic wake oscillations, in a flowing $e$-$i$ plasma.
@@
Nonlinear modelling: Theoretical analysis confirms simulation findings.
@@
Potential application: These new insights into plasma-debris interactions may be useful for management of space debris in low Earth orbit.
@@ 1 Introduction
In recent times, our understanding about plasma-debris interaction has been steadily progressing due to several theoretical, experimental, and numerical studies [18, 19, 20, 21, 22, 23]. It is now definitively proved both theoretically and through numerical simulations that the depending on the nature of the charged debris, different nonlinear waves in the ion-acoustic regime can get excited by moving debris [21, 23]. While we usually see dispersive shock waves due to a positively charged external debris, pinned solitons are formed due to phase space trapping through ion-ion counter streaming instability (IICSI) due to a negatively charged debris [23]. Some very recent laboratory and numerical simulation studies have looked into the plasma-debris interaction in multi-ion plasma [22]. Several other studies have focused on shape, dimension, and speed of the debris on the propagation of these solitons theoretically and with experiments [24, 20], as well as bending and acceleration [25], damping (both collisional and Landau) [26] of the excited structures.
@@
In this work, we plan to explore a yet unexplored area related to plasma-debris interaction by allowing the charge in the external debris to vary over time along with the presence of a stochastic component. There are good reasons as to why the charge on external debris should vary. Some seemingly plausible causes are periodic photoemission due to eclipse transitions (especially debris in the LEO), passage of the debris through the regions of density and/or temperature gradient of the ambient plasma, plasma wave interactions, intrinsic forcing due to spin or tumbling motion of the debris, and external modulations viz. during spacecraft remediation processes, among others. Out of these scenarios, the most important one is perhaps the debris charge modulation due to plasma wave interactions as this is an inherent process and possibly cannot be avoided. Debris moving at or near ion-acoustic speed is expected to make the charge on debris oscillate in time at ion-plasma frequency $(\omega_{pi})$, through ion-acoustic wake oscillations [27]. As an ion-acoustic wave (IAW) passes over the debris, it periodically alters the ion and electron densities. Since the ion and electron currents depend sensitively on these densities, the wave creates an imbalance in the ion and electron fluxes to the debris surface. This imbalance, in turn, disturbs the equilibrium floating potential of the debris and the debris charge adjusts dynamically to restore the balance, leading to a periodically oscillating debris charge. So, basically IAWs excited by the external charged debris, itself can cause the periodic fluctuation of the debris charge at frequencies $\sim\omega_{pi}$, which is very much inherent to the debris-plasma interaction scenario. This process is very much similar to the process of dust-charge fluctuation in a dusty plasma [27, 28]. This inherent charge fluctuation of the debris can be thought to be a forcing term, which can result in parametric coupling and can give rise to a wide range of nonlinear phenomena including chaos, sheath instability, nonlinear Landau damping [29] (NLLD) and can also affect the local turbulence. In this analysis, we primarily focus on two effects – chaos and NLLD, the signatures of which are most prominent. While detecting chaos is pretty straight forward, the same is not true for NLLD, which requires more than one confirmation methods. We believe that the periodic oscillation near $\omega_{pi}$ can trigger NLLD, which can then limit the wave amplitude, providing a feedback mechanism on the charging process and may eventually affect the nonlinear structures [30]. On the other hand, as chaotic oscillations can imply the breakdown of coherent wave-particle energy exchange, it can act as a probe for the diagnostic tool the wave-particle interaction limit [31, 32].
@@
The paper is organized as follows. In Section II, we present the results of kinetic and fluid simulations with the help of our h-PIC-MCC and mFCT codes, respectively. In Section III, we provide a theoretical analysis of the results obtained in Section II. We provide an additional analysis of the results obtained in earlier sections, in terms of power spectrum analysis in Section IV. In Section V, we conclude. We also provide a brief analysis of the effect of fluctuation of debris charge on turbulent dynamics.
@@ 2.1 The premise
Though technically it should be possible to model a self-consistent periodic fluctuation of charge in an external debris embedded in a plasma, in reality it is much harder to model it mathematically as it would naturally require calculation of collision cross-sections to exactly determine the electron and ion currents to the debris, which may not obey the much used orbit motion limited (OML) theory, usually used to calculate the currents. It may be further complicated by extended and arbitrary shape of the debris. Besides, a self-consistent fluctuation will necessarily require a coupling of a circuit-like differential equation with the partial differential equations of the fluid model for the plasma, so that we shall require a charging equation like
@@
In order to avoid these complexities, we have chosen to model the periodic fluctuation of charge on external debris through an “external forcing” equation for the debris charge density $\rho_{{\rm ext}}$
$\rho_{{\rm ext}}(t)=\rho_{0}{\cal F}(t),$ (2)
@@ 2.2 The plasma model
where the symbols have their usual meanings. The pressure for ions is treated as polytropic
@@
where $\gamma$ is the ratio of specific heats. The last term in the Poisson equation is the term due to external charged debris which is moving with respect to the bulk plasma at a constant velocity $v_{d}$. Note that depending on whether the debris charged positively or negatively, $\rho_{{\rm ext}}\lessgtr 0$. The normalized model can be written as
@@
where the densities are normalised by the equilibrium plasma density $n_{0}$, length is normalized by electron Debye length $\lambda_{D}$, time is normalized by the inverse of ion-plasma frequency $\omega_{pi}$, plasma potential $\phi$ is normalized by $(T_{e}/e)$. The temperature is expressed in energy units and and velocity is normalized by the ion-acoustic speed $c_{s}=\sqrt{T_{e}/m_{i}}$. The quantity $\sigma$ is the ratio of ion to electron temperature and is usually $\ll 1$. In the above equations, the external charge density $\rho_{{\rm ext}}$ is also normalized accordingly.
@@ 2.3 Kinetic model – PIC simulation
In this section, we shall carry out a PIC simulation of the above-mentioned $e$-$i$ plasma with a time-varying external charged debris. The simulation is carried out with our well-tested hybird-PIC-MCC code [33, 34, 35, 23], which has already been successfully applied to several electrostatic problems including plasmas with secondary electron emission, kinetic dust-ion-acoustic wave with ion-Landau damping effect, and $e$-$i$ dusty plasma with external charged debris.
@@
The plasma is assumed to have an $e$-$i$ number density $n_{e,i}\sim 10^{16}\,{\rm m}^{-3}$, with electron temperature $T_{e}\sim 1\,{\rm eV}$ and ion temperature $T_{i}\sim 0.01\,{\rm eV}$. The simulation box length $L\sim 140\lambda_{D}\sim 0.01\,{\rm m}$. The code can handle any arbitrary-shaped external charge distribution of arbitrary strength which can be either positive or negative with time-varying charge density. We can either make the debris move across the plasma or vice versa – the results being same in either case. Computationally, as mentioned before, we model our time-varying charge density with the following function
$\rho_{{\rm ext}}(t)=\rho_{0}[1+\tanh\{\Delta\sin(\nu t)\}]+\rho_{{\rm noise}}(t),$ (12)
@@
which is a symmetric (Gaussian-like) charge distribution with $\Delta$ as the flatness parameter of the time variation and $\nu$ as the oscillation frequency of the charge distribution. The larger the value of $\Delta$ is, the steeper is the variation. Note that the $\rho_{{\rm noise}}(t)$ part denotes a stochastic part of $\rho_{{\rm ext}}(t)$, which adds a degree of random-ness to the oscillation. The variable data $(v_{i},n_{i})$ are extracted across the simulation domain as independent time series, which are then examined with Wolf’s algorithm for possible chaotic nature. It is now well known that Wolf’s algorithm [36] applies Taken’s theorem [37] to construct an $m$-dimensional phase space from a scalar time series $\{x_{t}\}$,
@@
with $m_{{\rm PIC}}$ as the embedding dimension and $\tau$ as the time delay. The algorithm continuously measures the divergence between two nearest neighbor (forward in time)
@@
where $\bm{X}^{\prime}_{t}$ is $\bm{X}_{t}$’s nearest neighbor (in time). The so-called Lyapunov exponent [38], which is nothing but the exponential divergence of the nearest trajectories, is then calculated over valid divergence segments as
@@
where $\Delta t$ time difference taken to compute the divergence. The maximal Lyapunov exponent is then calculated by
@@
where $N$ is the number of valid divergence segments. Any positive $\lambda$ indicates chaos and the further $\lambda$ is away from 00, the stronger is the chaos. A typical rule of thumb is to have $\lambda\gtrsim 0.01$ for a definitive proof of chaos. In our case, we have used an embedding dimension $m_{{\rm PIC}}=4$ for our chaos test.
Figure 1: Maximal Lyapunov exponent $\lambda$, shown in the $x$-$\nu$ plane for a positively charged external debris (shown as the orange strip). The blue and red points indicate values of $\lambda$ with and without stochastic fluctuation. The results are at time $\sim 20\tau_{pi}$.
Figure 2: Maximal Lyapunov exponent $\lambda$, shown in the $x$-$\nu$ plane for a negatively charged external debris (shown as the orange strip). The blue and red points indicate values of $\lambda$ with and without stochastic fluctuation as before. The results are at time $\sim 20\tau_{pi}$.
The result of the simulation is shown in Figs.1 and 2, where we have shown the scatter plots of $\lambda$ in a $x$-$\nu$ plane. The position of the debris is indicated by an orange-colored strip in the panels. The position is expressed in terms of electron Debye length $\lambda_{D}$ and the frequency $\nu$ is expressed in terms of ion plasma frequency $\omega_{pi}$. In order to help the perspective visualization of the value of $\lambda$, we have added “vertical drop-lines” from every point to the $z=0$ baseline so that the length of the ‘drop-line’ indicates the value of $\lambda$. All simulation results are at time $\sim 20\tau_{pi}$, where $\tau_{pi}$ is the plasma period $=\omega_{pi}^{-1}$. The results for positively charged debris is shown in Fig.1 and the negatively charged results are shown in Fig.2. As we can see from Fig.1 that for positively charged debris, increase in debris velocity definitely induces quite stronger chaos in the oscillation, specifically at the site of the debris. The average strength of the $\lambda$ in panel (a) is almost in the boundary line of chaos $\sim 0.01$, which is without any relative motion between the plasma and external debris $(v_{d}=0)$. This should be contrasted with those in panels (b), (c), and (d), where we can see a definitive built up of chaotic oscillations. It should also be noted that the site of chaotic oscillation is mostly confined to the debris site. Besides, the stochasticity of fluctuation is not seen to have any systematic and correlated effect on $\lambda$, a situation which we shall explain later in Section 3.2.4.
@@
The corresponding ion phase space and ion density plots are shown in Fig.3 for positive debris and in Fig.4 for negative debris. All frames of Figs.3 and 4 are drawn in the rest frame of the debris, which lies in the middle of the simulation window denoted by a dashed vertical line in each panel. The ion velocity is normalized to the ion-acoustic speed and the ion density is normalized by its equilibrium value. The length is normalized by electron Debye length $\lambda_{D}$. All these plots are without any stochastic fluctuation component. What we can see from the plots is that, the periodic charge fluctuation of the external debris has a minimal effect when the debris charge is negative, also manifested by weak chaos, as shown before.
Figure 3: Ion phase space plots and ion density plots are shown in panels (a), (b) and density (c), (d) for positively charged external debris. The left hand plots (black colored) are without any fluctuation of the debris charge and the right hand plots (red colored) are for periodic fluctuation of the debris charge. The value of the fluctuation frequency $\nu$ is given in each panel.

### v3

@@ Abstract
Model: Investigates the effect of periodic charge fluctuation of external debris in a flowing $e$-$i$ plasma.
@@
Nonlinear modeling: Theoretical analysis confirms simulation findings.
@@
Potential application: Findings offer new insights into the plasma-debris interactions which may add another dimension to the debris detection scenario in low Earth orbit plasma.
@@ 1 Introduction
In recent times, our understanding about plasma-debris interaction has been steadily progressing due to several theoretical, experimental, and numerical studies [18, 19, 20, 21, 22, 23]. It is now definitively proved both theoretically and through numerical simulations that depending on the nature of the charged debris, different nonlinear waves in the ion-acoustic regime can get excited by moving debris [21, 23]. While we usually see dispersive shock waves due to a positively charged external debris, pinned solitons are formed due to phase space trapping through ion-ion counter streaming instability (IICSI) for a negatively charged debris [23]. Some very recent laboratory and numerical simulation studies have looked into the plasma-debris interaction in multi-ion plasma [22]. Several other studies have focused on shape, dimension, and speed of the debris on the propagation of these solitons theoretically and with experiments [24, 20], as well as bending and acceleration [25], damping (both collisional and Landau) [26] of the excited structures.
@@
In this work, we plan to explore a yet unexplored area related to plasma-debris interaction by allowing the charge in the external debris to vary over time along with the presence of a stochastic component. There are good reasons as to why the charge on external debris should vary. Some seemingly plausible causes are periodic photoemission due to eclipse transitions (especially debris in the LEO), passage of the debris through the regions of density and/or temperature gradient of the ambient plasma, plasma wave interactions, intrinsic forcing due to spin or tumbling motion of the debris, and external modulations viz. during spacecraft remediation processes, among others. Out of these scenarios, the most important one is perhaps the debris charge modulation due to plasma wave interactions as this is an inherent process and possibly cannot be avoided. Debris moving at or near ion-acoustic speed is expected to make the charge on debris oscillate in time at ion-plasma frequency $(\omega_{pi})$, through ion-acoustic wake oscillations [27]. As an ion-acoustic wave (IAW) passes over the debris, it periodically alters the ion and electron densities. Since the ion and electron currents depend sensitively on these densities, the wave creates an imbalance in the ion and electron fluxes to the debris surface. This imbalance, in turn, disturbs the equilibrium floating potential of the debris and the debris charge adjusts dynamically to restore the balance, leading to a periodically oscillating debris charge. So, basically IAWs excited by the external charged debris, itself can cause the periodic fluctuation of the debris charge at frequencies $\sim\omega_{pi}$, which is very much inherent to the debris-plasma interaction scenario. This process is very much similar to the process of dust-charge fluctuation in a dusty plasma [27, 28]. This inherent charge fluctuation of the debris can be thought to be a forcing term, which can result in parametric coupling and can give rise to a wide range of nonlinear phenomena including chaos, sheath instability, nonlinear Landau damping [29] (NLLD) and can also affect the local turbulence. In this analysis, we primarily focus on two effects – chaos and NLLD, the signatures of which are most prominent. While detecting chaos is pretty straight forward, the same is not true for NLLD, which requires more than one confirmation methods. We believe that the periodic oscillation near $\omega_{pi}$ can trigger NLLD, which can then limit the wave amplitude, providing a feedback mechanism on the charging process and may eventually affect the nonlinear structures [30]. On the other hand, as chaotic oscillations can imply the breakdown of coherent wave-particle energy exchange, it can act as a probe for the wave-particle interaction limit [31, 32]. We carry out our primary investigation with a 1D electrostatic particle-in-cell (PIC) method [33]. We shall also use the results of a flux-corrected-transport (FCT) method [34] to support our findings.
1.1 Why 1D?
While the real physical system is invariably 3D, one can get valuable physics information from 1D models very easily. There are three pressing reasons why in many cases, we resort to a 1D modeling. First, the symmetry. As there is no preferential direction, any direction is equivalent. Second, in many cases, the dominant dynamics happen only in one direction viz. direction of the flow. So, one can analyse the essential physics from studying the system in that dominant direction only. The third and the most important is the ease of calculation in 1D models as compared to 2D or 3D. In 2D or 3D, the resolution required to resolve Debye-scale fluctuations in the vicinity of the external debris will be prohibitively large.
In our case, as we shall see, the major dynamical activities are along the longitudinal direction or the direction of plasma (or debris) flow. So as long as we do not expect the transverse activities to affect the dynamics considerably (which is in our case), we can safely consider the 1D model. By assuming a slab-symmetric or axisymmetric configuration and a uniform plasma flow, the system effectively reduces to one dimension without significant loss of physical generality. Note that when an extended object (like a block or slab) move in a plasma, majority of the dynamics happen across the plasma-facing side of the debris, which is essentially like a 1D situation as all interactions across the plasma-facing side should be same. The situation would have been completely different had there been a magnetic field, which would have necessarily required transverse dynamics, thus forcing a multi-dimensional model.
Moreover, the 1D model allows us to use tractable mathematical techniques for analysing the chaos such as Lyapunov exponent analysis, Kolmogorov–Arnold–Moser (KAM) theory, and finite time Lyapunov exponent (FTLE) maps, which would be analytically and computationally prohibitive in higher dimensions. Hence, the 1D approach is not only physically motivated but also essential for uncovering the fundamental mechanisms governing plasma-debris interactions, which can be later benchmarked in detail through 2D or 3D models (which is beyond the scope of this work). The 1D model thereby serves as a minimal yet powerful framework to isolate and study these nonlinear mechanisms.
1.2 Organisation of the paper
@@
The paper is organized as follows. In Section II, we present the results of kinetic and fluid simulations with the help of our h-PIC-MCC and mFCT codes, respectively. In Section III, we provide a theoretical analysis of the results obtained in Section II. We provide an additional analysis of the results obtained in earlier sections, in terms of power spectrum analysis in Section IV. In Section V, we conclude. We also provide a brief analysis of the effect of fluctuation of debris charge on turbulent dynamics. In the Appendix, we discuss the validity of our kinetic model in the light of $e$-$i$ collisions.
@@ 2.1 The premise
Though technically it should be possible to model a self-consistent periodic fluctuation of charge in external debris embedded in a plasma, in reality it is much harder to model it mathematically as it would naturally require calculation of collision cross-sections to exactly determine the electron and ion currents to the debris, which may not obey the much used orbit motion limited (OML) theory, usually used to calculate the currents. It may be further complicated by the extended and arbitrary shape of the debris. Besides, a self-consistent fluctuation will necessarily require a coupling of a circuit-like differential equation with the partial differential equations of the fluid model for the plasma, so that we shall require a charging equation like
@@
In order to avoid these complexities, we have chosen to model the periodic fluctuation of the charge on external debris through an “external forcing” equation for the debris charge density $\rho_{{\rm ext}}$
$\rho_{{\rm ext}}(x,t)=\rho_{0}(x){\cal F}(t),$ (2)
@@
Our primary results are due to the PIC simulation as it incorporates all the kinetic effects and full nonlinearities and also model a realistic plasma situation. The FCT model is a reduced fluid model, results from which should largely support the PIC simulation results.
@@ 2.2 The plasma model
where $n_{i}$ and $n_{e}$ are the ion and electron number densities respectively, $v_{i}$ is the ion fluid velocity, $p_{i}$ is the ion pressure, $\phi$ is the electrostatic potential, $m_{i}$ is the electron mass, $e$ is the electronic charge, $\epsilon_{0}$ is the vacuum permittivity and $T_{e}$ is the electron temperature. The last term in the Poisson equation, $\rho_{ext}$ is the term due to external charged debris which is moving with respect to the bulk plasma at a constant velocity $v_{d}$. Note that depending on whether the debris is charged positively or negatively, $\rho_{{\rm ext}}\lessgtr 0$. The pressure for ions is treated as polytropic
@@
where $\gamma$ is the ratio of specific heats. The normalized model can be written as
@@
where the densities are normalised by the equilibrium plasma density $n_{0}$, length is normalized by electron Debye length $\lambda_{D}$, time is normalized by the inverse of ion-plasma frequency $\omega_{pi}$, plasma potential $\phi$ is normalized by $(T_{e}/e)$. The temperature is expressed in energy units and velocity is normalized by the ion-acoustic speed $c_{s}=\sqrt{T_{e}/m_{i}}$. The quantity $\sigma$ is the ratio of ion to electron temperature and is usually $\ll 1$. In the above equations, the external charge density $\rho_{{\rm ext}}$ is also normalized accordingly.
@@ 2.3 Kinetic model – PIC simulation
2.3.1 The PIC simulation
It is important for the general reader that we explain the basics of PIC simulation of plasma, though it is widely used and described in details in many good textbooks, especially one by Birdsall [33]. The dynamics modelled by PIC simulation is as per the Boltzmann-Vlasov (or simply Vlasov) equation [35]
$\frac{\partial f_{j}}{\partial t}+\bm{u}_{j}\cdot\nabla f_{j}+\frac{q_{j}}{m_{j}}\bm{E}\cdot\nabla_{u_{j}}f_{j}=0,$ (12)
where $j=i,e$ stands for ions and electrons, $f_{j}$ are the respective velocity distributions, $\bm{u}_{j}$ are the respective particle velocities, in contrast to fluid velocity $v_{i}$ used in Eqs.(8-10), and $\bm{E}$ is the plasma electric field. The operator $\nabla_{u_{j}}$ is the gradient operator in the respective velocity space. In presence of non-zero collisional momentum exchange, the right hand side of the above equation must be replaced by a collision term (which is zero in our case). The fluid equations given by Eqs.(8-10) are nothing but differently weighted average of the Vlasov equation Eq.(12). While the Vlasov equation is fully kinetic, the fluid model looses out the kinetic effects (those effects which involves individual particle properties) due to the averaging. Naturally, the fluid model is an approximation to the kinetic model.
The crux of the PIC model lies in its ability to model a large number of real-life plasma particles (of each kind) through a computational particle (also known as so-called super-particle) [33], which greatly reduces the computation time. Unless we consider a very strongly coupled plasma, the PIC model proves to be a very efficient method to model kinetic plasma phenomena [33]. The PIC model starts with distribution of the computational particles as per the given velocity distribution (which in our case, is a Maxwellian) on a grid, which is a 1D uniform grid in our case. Our particle distribution is spatially random and uniform. Once the particles are distributed, the electrostatic potential $\phi$ is calculated at the grid points through a nearest neighbour approximation [33, 36], which is then used to calculate the plasma electric field $\bm{E}$ at the grid points through the relation
$\bm{E}=-\nabla\phi.$ (13)
This provides us with the force $\bm{F}$ at the particle positions through $\bm{F}=q\bm{E}$, which is again interpolated through a nearest neighbour scheme [33, 36]. Once the force is obtained, the following equations are inverted subsequently to update new velocity $\bm{u}$ and then new position $\bm{x}$,
$\frac{d\bm{u}}{dt}$ $=$ $\frac{\bm{F}}{m},$ (14)
$\frac{d\bm{x}}{dt}$ $=$ $\bm{u}$ (15)
and the cycle repeats.
2.4 Description of the simulation
@@
In this section, we shall present the results of PIC simulation of the above-mentioned $e$-$i$ plasma with a time-varying external charged debris. The simulation is carried out with our well-tested hybird-PIC-MCC code [36, 37, 38, 23], which has already been successfully applied to several electrostatic problems including plasmas with secondary electron emission, kinetic dust-ion-acoustic wave with ion-Landau damping effect, and $e$-$i$ dusty plasma with external charged debris.
@@
The plasma is assumed to have an $e$-$i$ number density $n_{e,i}\sim 10^{16}\,{\rm m}^{-3}$, with electron temperature $T_{e}\sim 1\,{\rm eV}$ and ion temperature $T_{i}\sim 0.01\,{\rm eV}$. The simulation box is of length $L\sim 140\lambda_{D}\sim 0.01\,{\rm m}$. The cell size for computation is $\sim 6.25\times 10^{-6}$ m which is equivalent to $\sim 0.05\lambda_{D}$, enough to resolve fine-scale, Debye-scale fluctuations. The simulation is being performed with periodic boundary conditions. The simulation time and the box length is adjusted, so that the nonlinearities in the neighbourhood of the debris are not affected by the boundaries. The code can handle any arbitrary-shaped external charge distribution of arbitrary strength which can be either positive or negative with time-varying charge density. We can either make the debris move across the plasma or vice versa – the results being same in either case. Computationally, as mentioned before, we model our time-varying charge density with the following function
$\rho_{{\rm ext}}(x,t)$ $=$ $\rho_{0}(x)[1+\tanh\{\Delta\sin(\nu t)\}]+\rho_{{\rm noise}}(t),$ (16)
$\rho_{0}(x)$ $=$ $\tilde{\rho}_{0}e^{-x^{2}/w},$ (17)
which is a spatially symmetric (Gaussian-like) charge distribution $\rho_{0}(x)$, oscillating in time. The parameter $\Delta$ is the flatness parameter of the time variation, $\nu$ is the oscillation frequency of the charge distribution, and $w$ is the measure of spatial width of the distribution which is $\sim 2\lambda_{D}$, with $\tilde{\rho}_{0}$ denoting the peak of the debris charge density. It should be noted that the value of $\tilde{\rho}_{0}$ must be more than a critical value in order to sustain the debris-induced nonlinearity. In our case, with the present plasma parameters, the minimum $\tilde{\rho}_{0}^{\textrm{(min)}}\sim 0.08$ with respect to the background charge density [23]. If it is not enough, the debris will be overwhelmed by the inherent noise in the system and the debris-induced effects will diffuse within a few plasma periods. The larger the value of $\Delta$ is, the steeper is the variation. Note that the $\rho_{{\rm noise}}(t)$ part denotes a stochastic part of $\rho_{{\rm ext}}(x,t)$, which adds a degree of random-ness to the oscillation.
Figure 1: A schematic (not to scale) representation of the spatial distribution $\rho_{0}(x)$ (left) for the external charge and temporal variation of the distribution (right). These representations are without stochastic noise.
2.4.1 Charge modulation
In Fig.1, we have schematically shown the spatial distribution of the external charge and its temporal variation (without noise). While the spatial distribution is a symmetric Gaussian, the distribution oscillates much like a top-hat function, oscillating between zero and some positive (or negative) value.
Or in other words, the charge distribution always remain positive or negative definite during the oscillation. The primary reason for choosing the oscillation is the way external debris get charged in the plasma. If we do not alter the debris properties (or plasma properties), the debris can get charged only in one way – either positive or negative, not both. One can, however get external debris charged in both ways if the mechanism of charging becomes completely different over time like the lunar surface. We know that the Moon acts like an external entity in the flow of solar wind plasma. During day-time, the lunar surface gets charged positively due to the solar UV radiation inducing photoemission of electrons from the lunar surface, while during night-time, the surface gets negatively charged. But, in our case, this fluctuation of charge on the debris is due to inherent plasma oscillations where there is no way that the charging mechanism can get changed within the oscillation time.
If we go a bit deeper, we see that the whole process of charging of external entities in a plasma is due to the asymmetric response of plasma electrons and positive charges toward the external surface. So, we can say that the top-hat-like fluctuation is rooted in the asymmetric response of the plasma.
2.4.2 Modeling of the noise
We treat the noise as one which can naturally arise in a plasma due to random fluctuations. If we denote the driving signal as $S(t)$, which in our case may be a periodic signal due to plasma wave, the total signal $T(t)$ can be thought to be composed of the signal $S(t)$ and a stochastic noise $R(t)$. However, we assume that the stochastic fluctuation is purely environmental i.e. it depends on the signal $S(t)$, so that the total signal can be written as
$T(t)=S(t)+R(t),$ (18)
where
$R(t)=\begin{cases}R_{0}(t),&\textrm{if}\,S(t)=0,\\ |\delta R(t)|S(t),&\textrm{if}\,S(t)>0,\end{cases}\quad\delta R(t)<0.$ (19)
This model can be justified by considering the fact that even in the absence of any driving signal, the plasma continues to have background thermal fluctuations such as spontaneous Langmuir oscillations [39]. This explains the relation
$T(t)=R_{0}(t),\quad\textrm{if}\,S(t)=0.$ (20)
On the other hand, when there is a driving signal (such as in our case), the plasma responds nonlinearly, which can be through nonlinear wave-particle interactions (viz. nonlinear Landau damping) [40, 41]. As this response is supposed to be dependent on the driving signal, we have
$T(t)=S(t)[1-|\delta R(t)|].$ (21)
Regarding the strength of the stochastic component, we have assumed it to be $R_{0}(t)\leq S(t)$ and $|\delta R(t)|\leq 1$. Though there is not any hard justification to this assumption, it is quite intuitive that stochastic fluctuations are always supposed to be lower than the driving signal, when the driving signal goes to zero and it always attenuates the signal, whenever it is non-zero [40, 41].
2.4.3 Chaotic dynamics
We are now going to discuss the chaotic dynamics induced by different oscillating external charge distributions.
@@ 2.4.3 Chaotic dynamics
The variable data $(v_{i},n_{i})$ are extracted across the simulation domain as independent time series, which are then examined with Wolf’s algorithm for possible chaotic nature. It is now well known that the Wolf’s algorithm [42] applies Taken’s theorem [43] to construct an $m$-dimensional phase space from a scalar time series $\{x_{t}\}$,
@@
with $m_{{\rm PIC}}$ as the embedding dimension and $\tau$ as the time delay. The algorithm continuously measures the divergence between two nearest neighbour (forward in time)
@@
where $\bm{X}^{\prime}_{t}$ is $\bm{X}_{t}$’s nearest neighbour (in time). The so-called Lyapunov exponent [44], which is nothing but the exponential divergence of the nearest trajectories, is then calculated over valid divergence segments as
@@
where $\Delta t$ is the time difference taken to compute the divergence. The maximal Lyapunov exponent is then calculated by
@@
where $N$ is the number of valid divergence segments. Any positive $\lambda$ indicates chaos and the further $\lambda$ is away from 0, the stronger is the chaos. A typical rule of thumb is to have $\lambda\gtrsim 0.01$ for a definitive proof of chaos. In our case, we have used the variable $n_{i}$ for the scalar time series $\{x_{i}\}$ and an embedding dimension $m_{{\rm PIC}}=4$ for our chaos test.
Figure 2: Maximal Lyapunov exponent $\lambda$, shown in the $x$-$\nu$ plane for a positively charged external debris (shown as the orange strip). The blue and red points indicate values of $\lambda$ with and without stochastic fluctuation. The results are at time $\sim 20\tau_{pi}$. The arrow indicates the direction of debris velocity.
Figure 3: Maximal Lyapunov exponent $\lambda$, shown in the $x$-$\nu$ plane for a negatively charged external debris (shown as the orange strip). The blue and red points indicate values of $\lambda$ with and without stochastic fluctuation as before. The results are at time $\sim 20\tau_{pi}$. The arrow indicates the direction of debris velocity.
The result of the simulation is shown in Figs.2 and 3, where we have shown the scatter plots of $\lambda$ in a $x$-$\nu$ plane. The position of the debris is indicated by an orange-colored strip in the panels. The position is expressed in terms of electron Debye length $\lambda_{D}$ and the frequency $\nu$ is expressed in terms of ion plasma frequency $\omega_{pi}$. In order to help the perspective visualization of the value of $\lambda$, we have added “vertical drop-lines” from every point to the $z=0$ baseline so that the length of the ‘drop-line’ indicates the value of $\lambda$. All simulation results are at time $\sim 20\tau_{pi}$, where $\tau_{pi}$ is the plasma period $\omega_{pi}^{-1}$. The results for positively charged debris is shown in Fig.2 and that for negatively charged debris are shown in Fig.3. As we can see from Fig.2 that for positively charged debris, increase in debris velocity definitely induces quite stronger chaos in the oscillation, specifically at the site of the debris. The average strength of the $\lambda$ in panel (a) is almost in the boundary line of chaos $\sim 0.01$, which is without any relative motion between the plasma and external debris $(v_{d}=0)$. This should be contrasted with those in panels (b), (c), and (d), where we can see a definitive built up of chaotic oscillations. It should also be noted that the site of chaotic oscillation is mostly confined to the debris site. Besides, the stochasticity of fluctuation is not seen to have any systematic and correlated effect on $\lambda$, a situation which we shall explain later in Section 3.2.4.
@@
The corresponding ion phase space and ion density plots are shown in Fig.4 for positive debris and in Fig.5 for negative debris. All frames of Figs.4 and 5 are drawn in the rest frame of the debris, which lies in the middle of the simulation window denoted by a dashed vertical line in each panel. The ion velocity is normalized to the ion-acoustic speed and the ion density is normalized by its equilibrium value. The length is normalized by electron Debye length $\lambda_{D}$. All these plots are without any stochastic fluctuation component. What we can see from the plots is that, the periodic charge fluctuation of the external debris has a minimal effect when the debris charge is negative, also manifested by weak chaos, as shown before.
Figure 4: Ion phase space plots and ion density plots are shown in panels (a), (b) and (c), (d) for positively charged external debris. The left hand plots (black colored) are without any fluctuation of the debris charge and the right hand plots (red colored) are for periodic fluctuation of the debris charge. The value of the fluctuation frequency $\nu$ is given in each panel.